5 ltr of an ieal gas initially at 1 atmosphere pressure is compressed adiabatically to 1 lt.Calculate the final pressure[Specific heat capacities of the gas at constant pressure and at constant volume are 525 J/kg k and 315J/kg k respectively]
 

Grade:11

1 Answers

Shiv Sagar
4 years ago
For an adiabatic process, P x V^(gamma) remains a constant. Here gamma is the ratio of Cp and Cv.
 
gamma = 525/315 = 1.66
 
P1 x V1^(1.66) = P2 x V2^(1.66)
1 x 5^1.66 = P2 x 1^1.66
Hence P2 = 5^1.66 = 14.62